### v3

# Ralph Abbey1, Jorge Silva1, Emiliano De Cristofaro3 1SAS   2UCL   3UC Riverside

Abstract
The Synthetic Minority Over-sampling Technique (SMOTE) is one of the most widely used methods for addressing class imbalance and generating synthetic data. Despite its popularity, little attention has been paid to its privacy implications; yet, it is used in the wild in many privacy-sensitive applications. In this work, we conduct the first systematic study of privacy leakage in SMOTE: we begin by showing that prevailing evaluation practices, i.e., naive distinguishing and distance-to-closest-record metrics, completely fail to detect any leakage and that membership inference attacks (MIAs) can be instantiated with high accuracy. Then, by exploiting SMOTE’s geometric properties, we build two novel attacks with very limited assumptions: DistinSMOTE, which perfectly distinguishes real from synthetic records in augmented datasets, and ReconSMOTE, which reconstructs real minority records from synthetic datasets with perfect precision and recall approaching one under realistic imbalance ratios. We also provide theoretical guarantees for both attacks. Experiments on eight standard imbalanced datasets confirm the practicality and effectiveness of these attacks. Overall, our work reveals that SMOTE is inherently non-private and disproportionately exposes minority records, highlighting the need to reconsider its use in privacy-sensitive applications and as a baseline for assessing the privacy of modern generative models.
1 Introduction
From rare disease diagnosis to fraud detection, machine learning tasks can be profoundly affected by severe class imbalance, where instances of interest – the minority class – are much rarer than the majority class [29]. Models often underperform under these conditions, exhibiting biases toward the majority and failing to capture the minority reliably [9]. One of the most influential and widely adopted approaches to address this is the Synthetic Minority Over-sampling Technique (SMOTE) [7], which augments the imbalanced data by upsampling or generating synthetic samples of the underrepresented class through linear interpolation between minority records. Due to its simplicity and effectiveness, SMOTE continues to play a central role in real-world applications. To put things in context, the SMOTE paper has been cited nearly 40k times, Microsoft Azure offers built-in SMOTE components [43, 44], and most MLaaS services support it [26, 5]. Overall, SMOTE is primarily used in two contexts: 1) as a data augmentation technique for machine learning classifiers, and 2) as a synthetic data generation method to facilitate data sharing.
Data Augmentation. SMOTE was originally proposed as a pre-processing/upsampling technique to augment the real dataset, thus improving classifier performance, especially F1 score and recall, when trained on the augmented data. Practitioners rely on SMOTE in a wide range of medical applications, including cancer diagnosis [21], heart-related diseases [46, 20], diabetes prediction [54, 2], genetic risk prediction [36], etc. Beyond medicine, SMOTE is widely applied in finance, particularly for credit-card fraud detection [64, 34] and predicting customer churn [52, 49].
Synthetic Data. SMOTE has also gained traction as a method for generating synthetic tabular data. Often used as a baseline for more advanced models like GANs and VAE, SMOTE has been shown to perform on par with, or even better than, generative approaches [42, 35]. Moreover, its extensive use as a baseline has led to modern diffusion-based models (not explicitly designed with privacy in mind) to be characterized as privacy-preserving simply because they outperform SMOTE, a pattern repeatedly observed in top-tier machine learning publications [37, 63, 50, 45]. SMOTE is also applied for medical synthetic data [33], and beyond machine learning research, it has been recognized as a promising technique for improving access to census data by public sector entities [48, 1].
Figure 1: DistinSMOTE and ReconSMOTE attacks vs. augmented/synthetic data generated by SMOTE.
Augmented Data
Naive MIA DistinSMOTE
0.01 $\pm$ 0.01 0.68 $\pm$ 0.07 1.00 $\pm$ 0.00
Synthetic Data
Naive MIA ReconSMOTE
0.16 $\pm$ 0.10 0.93 $\pm$ 0.02 1.00 $\pm$ 0.00
Table 1: Performance of privacy attacks vs. SMOTE. Naive refers to the current privacy evaluation practices. MIAs are applied to SMOTE for the first time. DistinSMOTE and ReconSMOTE are our novel attacks.
Roadmap. Although not originally designed with privacy in mind, SMOTE is extensively used in sensitive public-facing applications that process personal data. However, its privacy risks have often been overlooked or significantly underestimated. In this paper, we fill this gap by studying whether, why, and how much privacy leakage occurs from using SMOTE either as a data augmentation method or as a standalone synthetic data generator.
We begin by showing that SMOTE appears to have no privacy leakage when evaluating it through the current practice of training a classifier to distinguish real from synthetic records in augmentation settings or that of measuring the distance from synthetic to real records (more precisely, the distance to closest record, or DCR [65]). We refer to the former as naive distinguishing and the latter as naive metrics. We also instantiate—to our knowledge, for the first time—a Membership Inference Attack (MIA) [56, 59] against SMOTE, showing that attackers can accurately infer whether a target record was part of the real training data.
Next, we propose two novel near-perfect privacy attacks with minimal and realistic assumptions: a Distinguishing (DistinSMOTE) and a Reconstruction attack (ReconSMOTE). Both only assume access to a single augmented or synthetic dataset and knowledge that SMOTE generated it (see Figure 1). By exploiting SMOTE’s geometric properties, DistinSMOTE distinguishes real from synthetic records in augmentation settings, while the more ambitious ReconSMOTE reconstructs real minority records from synthetic data. We also provide a theoretical analysis for both attacks, showing they run at worst in $\mathcal{O}(n^{2}d+n(kr)^{2})$, where $n$, $d$, and $k$ denote, respectively, the number of input records, features, and SMOTE neighbors, and $r$ represents the data imbalance ratio. While quadratic in $n$, the complexity remains practical (especially with optimized search), with both attacks running within minutes on all datasets we experiment with.
DistinSMOTE achieves perfect precision and recall, while ReconSMOTE reaches perfect precision – which is more critical in privacy attacks [6] – with recall increasing exponentially (with rate $\approx{r}/{k}$), reaching 1 under realistic parameter values ($k=5$, $r\geq 20$).
Our experiments, summarized in Table 1, on eight standard imbalanced datasets demonstrate that:
  • Naive distinguish (0.01 precision)/metrics (0.16 accuracy) completely underestimate risks.
  • State-of-the-art MIAs achieve 0.68 AUC on augmented and 0.93 on synthetic data for 100 vulnerable targets, although being time-consuming. Also, sensitivity of targets increases when classifiers are trained on augmented vs. real data, yielding a 17% rise in MIA AUC.
  • DistinSMOTE perfectly detects the real records in an augmented dataset.
  • ReconSMOTE achieves perfect precision when reconstructing real minority records from a single synthetic dataset. While its average recall is 0.85, it reaches 1 for imbalance ratios of 20 or higher, consistent with our theoretical analysis.
Implications. Our findings provide further evidence that privacy cannot be treated as an afterthought when applying non-private techniques like SMOTE in sensitive settings. Its use not only risks exposing individual records but can also undermine trust in data-driven systems that rely on synthetic data. Overall, our work has the following real-world implications for researchers/practitioners:
  1. 1.
    SMOTE is fundamentally non-private: its interpolation process makes privacy leakage inherent, not a matter of flawed implementation.
  2. 2.
    Minority records are disproportionately at risk: the very samples SMOTE aims to amplify and make more representative are also the most exposed.
  3. 3.
    SMOTE and DCR are unreliable: evaluating SMOTE with privacy metrics like DCR gives a misleading assessment of its privacy and should not be used to validate other generative models.
  4. 4.
    Caution is critical: performance gains from oversampling must be weighed vs. privacy risks.
2 Preliminaries
Notation. Let $D_{real}=(X,y)$ be a training dataset, where $X\subseteq\mathbb{R}^{n\times d}$ is the feature matrix consisting of $n$ samples with $d$-dimensional feature vectors and $y\in\{0,1\}^{n}$ the corresponding binary labels. The dataset consists of $n_{0}$ majority and $n_{1}$ minority records, with imbalance ratio $r=\tfrac{n_{0}}{n_{1}}>1$. In practice, $n_{0}\gg n_{1}$, which makes learning directly from the minority class challenging.
Algorithm 1 SMOTE [7]
1Real dataset $D_{real}$
2Number of neighbors $k$
3Augmented data $D_{aug}=D_{real}\cup D_{syn}$,
4  or Synthetic data $D_{syn}$
5Filter minority $X^{1}_{real}{\leftarrow}\{X_{real}[i]{\mid}y_{real}[i]{=}1\}$
6Compute $n_{1}=|X^{1}_{real}|$, $n_{0}=|D_{real}|-n_{1}$
7while $|D_{syn}|<n_{0}-n_{1}$ do
8  Randomly pick $x_{i}\in X^{1}_{real}$
9  Find $k$ nearest neighbors of $x_{i}$, $N(x_{i})$
10  Randomly choose $x_{j}\in N(x_{i})$
11  Sample $u\sim U(0,1)$
12  Generate $x_{syn}\leftarrow x_{i}+u(x_{j}-x_{i})$
13  Add $(x_{syn},1)$ to $D_{syn}$
14end while
SMOTE [7] addresses class imbalance by generating synthetic minority samples; see Algorithm 1. To create a synthetic record, a random minority record from $D_{real}$ is selected, one of its $k$ nearest minority neighbors is chosen, and a new point is generated by interpolating along the line segment between them. Repeating this process yields $D_{syn}$ with $n_{0}{-}n_{1}$ new samples, balancing the class distribution. The synthetic data can be used as a standalone synthetic dataset ($D_{syn}$) or to form an augmented dataset $D_{aug}=D_{real}\cup D_{syn}$, e.g., to improve classification performance.
Privacy Attacks. Membership Inference Attacks (MIAs) [56, 59] and Reconstruction Attacks [14, 3] are standard tools to empirically measure privacy leakage in ML. In MIAs, the adversary aims to infer whether a target record $(x_{T},y_{T})$ was part of the training dataset $D_{real}$. The attack can be framed as a repeated binary classification game: the adversary is given either a classifier (or a synthetic dataset) trained on $D_{real}$, or one trained on the neighboring $D_{real}^{\prime}=D_{real}\setminus{(x_{T},y_{T})}$, and infers which dataset was used. To do so, the adversary typically exploits differences in model behavior – such as prediction confidences on the target, or statistical features extracted from synthetic data.
In a reconstruction attack, the adversary aims to recover any full real records (i.e., an untargeted attack) from access to a released model or synthetic data. These attacks often assume access to auxiliary information, such as public data, accurate statistics, or limited query access to $D_{real}$.
We also consider distinguishing attacks, which are somewhat related to MIAs but focus on whether a record comes from the population-level data distribution rather than from the specific dataset used to train the model. In the context of SMOTE, we use these attacks to distinguish unlabelled records in $D_{aug}$ as either real ($D_{real}$) or synthetic ($D_{syn}$), since $D_{real}\cap D_{syn}=\emptyset$.
3 Related Work
As discussed in Section 1, SMOTE is widely used for data augmentation and synthetic data generation across various domains. Despite its popularity, prior work has focused primarily on its utility, while its privacy risks remain largely unexplored.
SMOTE has recently served as a baseline in several diffusion-based generative models [37, 63, 50, 45], all published at top-tier machine learning venues. A common pattern across these studies is the “SMOTE + DCR” workflow: they rely on the Distance to Closest Record (DCR) [65] as the primary privacy proxy, consistently reporting smaller DCR values for SMOTE and interpreting this as evidence that the newly proposed models are privacy-preserving. Kotelnikov et al. [37] additionally employ the “full black-box” attack from [8], which still reduces to DCR as its core signal.
More recently, this view has been challenged by Sidorenko and Tiwald [57], who show that some diffusion models [37, 45] actually achieve lower DCR values than SMOTE, thereby leaking more information about the training data. However, DCR itself has been shown to be an inadequate privacy metric – it consistently underestimates leakage [31, 4, 22] and does not correlate with leakage detected by MIAs [61]. To the best of our knowledge, despite its prominent role as a baseline, SMOTE has not yet been systematically evaluated with state-of-the-art MIAs or any model-specific attacks. This leaves a critical gap in understanding SMOTE’s true privacy risks and calls into question the validity of privacy claims across a recent line of generative-model research.
4 Privacy Attacks vs. SMOTE
In this section, we present our two novel privacy attacks that expose privacy leakage from SMOTE.
4.1 Adversarial Model
Assumptions. For both attacks, we assume an adversary with access to a single dataset generated by SMOTE ($D_{aug}$ for DistinSMOTE and $D_{syn}$ for ReconSMOTE). The adversary knows that the original SMOTE algorithm [7] was applied (Algorithm 1) and is aware of its parameters, specifically, the number of neighbors $k$ and the real data imbalance ratio $r$ (in practice, these can be approximated from the released dataset). The adversary relies solely on the geometrical properties of SMOTE to achieve its goals – either distinguishing or reconstruction. Unlike prior privacy attacks, no further knowledge is required: e.g., the adversary does not need access to public/representative data, repeated inference or generation, the model parameters, numerous shadow models or a meta-classifier (as in MIAs [59, 31, 4]), or published (accurate) aggregate statistics (as in reconstruction [14, 12]).
Objectives. For DistinSMOTE, the adversary aims to distinguish the real minority records from synthetic ones from observing $D_{aug}$, whereas for ReconSMOTE, to reconstruct them from $D_{syn}$. We focus on minority records from underrepresented regions of the feature space because they often correspond to the most vulnerable individuals. Such records carry the greatest privacy risks: they are easier to single out, more likely to be re-identified, and any disclosure disproportionately affects the individuals they represent [38, 59]. Indeed, regulators, including the UK Information Commissioner’s Office [32], have explicitly stressed the need to protect minority and outlier records.
We measure the attacks success using precision (the fraction of identified records that are truly real) and recall (the fraction of successfully identified real records), two standard metrics that together give a comprehensive view of performance. In privacy attacks, precision is especially critical, since even a handful of correctly identified records with high confidence can constitute a serious breach [6]. While secondary, capturing a large fraction of minority records is also important.
Data Assumptions. Our theoretical analysis of DistinSMOTE and ReconSMOTE relies on the following three realistic assumptions about the feature structure of the real minority records ($X^{1}_{real}$) and $k$:
Assumption 1 (Real-valued features).
All features in $X^{1}_{real}$ are real-valued.
Assumption 2 (Global non-collinearity).
No three distinct feature vectors in $X^{1}_{real}$ are collinear.
Assumption 3 (Minimum $k$).
The number of neighbors, $k\geq 3$.
In other words, all features are continuous and no three points lie on the same line. These assumptions align naturally with (high-dimensional) continuous data, are non-restrictive in practice, and, crucially, are satisfied by all datasets in our main experiments (see Section 5). Moreover, assuming $k\geq 3$ is necessary to uniquely identify the intersection point of the lines formed by vectors connecting neighboring points, and is a standard, practical choice in typical SMOTE configurations (default is 5).
4.2 Distinguishing Attack
In Algorithm 2, we outline the DistinSMOTE attack, which distinguishes real minority records from synthetic ones within an augmented dataset. The attack exploits the fact that, among any three collinear points, the middle one must be synthetic, since real points are non-collinear and SMOTE generates points strictly between them. The algorithm begins by searching from the convex hull of the minority records and iteratively explores neighbors inwards. When a collinear triplet is found, its midpoint is pruned from the candidate set with real points, and its neighbors are added to the queue for further inspection.
Algorithm 2 DistinSMOTE
1Augmented data $D_{aug}$
2Number of neighbors $k$, imbalance ratio $r$
3Detected real minority records $C^{1}$
4Filter minority $X^{1}_{aug}{\leftarrow}\{X_{aug}[i]{\mid}y_{aug}[i]{=}1\}$
5Initialize candidate set $C^{1}\leftarrow X^{1}_{aug}$
6Initialize queue $queue\leftarrow H(X^{1}_{aug})$ $\triangleright$ convex hull
7Initialize visited set $V\leftarrow\emptyset$
8while $queue\neq\emptyset$ do
9  for record $x_{i}\in queue$ do
10  if $x_{i}\notin V$ and $x_{i}\in C^{1}$ then
11  Add $x_{i}$ to $V$
12  Find $2\cdot k\cdot r$ nearest neighbors of $x_{i}$, $N(x_{i})$
13  for pairs of neighbors $(x_{j},x_{k})\in N(x_{i})$ do
14    if $x_{i},x_{j},x_{k}$ are collinear then
15    Identify middle point $x_{m}\in\{x_{i},x_{j},x_{k}\}$
16    Remove $x_{m}$ from $C^{1}$; Add $x_{m}$ to $V$
17    Add $N(x_{m})\cap C^{1}$ to queuequeue
18    end if
19  end for
20  end if
21  end for
22end while
23return $C^{1}$
Complexity. The nearest-neighbor search is the main cost. With brute-force search ($\mathcal{O}(nd)$ per query), each of the $n$ records requires finding $kr$ neighbors ($\mathcal{O}(nd)$) and checking all neighbor pairs ($\mathcal{O}((kr)^{2})$), yielding a worst-case complexity of $\mathcal{O}(n^{2}d+n(kr)^{2})$. In practice, optimized methods (e.g., KD/ball trees) reduce search to $\mathcal{O}(\log n)$ for small $d$ ($d\leq 32$ in our main datasets). Also, since $k$ and $r$ are typically (small) constants, the effective complexity is much lower, and the algorithm completes in under three minutes on all main datasets.
Accuracy Analysis. We analyze DistinSMOTE, with the theorem below formalizing the theoretical correctness of its labeling rule, achieving perfect precision and recall.
Theorem 1 (DistinSMOTE perfect precision & recall).
Under Assumptions 12, the labeling rule in the DistinSMOTE attack achieves perfect precision and recall (with probability 1).
Proof.
By SMOTE construction, each synthetic point lies strictly on the line segment between two real points ($x_{syn}=x_{i}+u(x_{j}-x_{i})$). Under the global non-collinearity assumption, no three real points are collinear, so any line in $D_{aug}$ containing at least three points consists of exactly two real endpoints $x_{i},x_{j}\in X^{1}_{real}$ and one or more synthetic interior points $x_{m_{1}},x_{m_{2}},\dots\in X^{1}_{syn}$.
DistinSMOTE follows this labeling rule: it finds lines via local search (lines 9-11 of Algorithm 2) and marks interior points as synthetic and endpoints as real (lines 12-13), while any real point not on such a line (i.e., not used in interpolation) is also labeled real. The local search guarantees all points are visited efficiently, and the labeling rule ensures all synthetic points are removed and all real points preserved. This leads to perfect precision and recall (with probability 1, except for negligible numerical precision effects that do not occur in practice). ∎
4.3 Reconstruction Attack
Algorithm 3 presents the ReconSMOTE attack, which operates solely on synthetic data. The attack relies on two main intuitions: 1) synthetic records lie along line segments connecting real minority points, so these lines can be detected by finding three or more collinear samples, and 2) such lines intersect exactly at the original real points. The algorithm begins by iteratively defining lines, searching each point and two of its neighbors for collinear triplets, and then extending them with additional collinear neighbors. For each line, the mean of its points is stored as a midpoint, providing a compact representation of the line’s location. Next, the algorithm examines pairs of midpoints to identify intersection points of the corresponding lines, which serve as candidate real records. Finally, we retain only intersections supported by at least three distinct lines, filtering out spurious candidates.
Algorithm 3 ReconSMOTE
1Synthetic data $D_{syn}$
2Number of neighbors $k$, imbalance ratio $r$
3Reconstructed real minority records $R^{1}$
4Filter minority $X^{1}_{syn}{\leftarrow}\{X_{syn}[i]{\mid}y_{syn}[i]{=}1\}$
5Initialize reconstructed set $R^{1}\leftarrow\emptyset$ and
6  line support map $S\leftarrow\emptyset$
7Initialize set of lines $\mathcal{L}\leftarrow\emptyset$, midpoints $\mathcal{M}\leftarrow\emptyset$
8Initialize visited set $V\leftarrow\emptyset$
9for record $x_{i}\in X^{1}_{syn}$ do
10  if $x_{i}\notin V$ then
11  Add $x_{i}$ to $V$
12  Find $2\cdot k\cdot r$ nearest neighbors of $x_{i}$, $N(x_{i})$
13  for pairs of neighbors $(x_{j},x_{k})\in N(x_{i})$ do
14  if $x_{i},x_{j},x_{k}$ are collinear then
15   Form initial line $(x_{i},x_{j},x_{k})$; Add $x_{j},x_{k}$ to $V$
16   for neighbor $x_{n}\in N(x_{i})\setminus\{x_{i},x_{j},x_{k}\}$ do
17   if $x_{n}$ collinear with line $(x_{i},x_{j},x_{k})$ then
18   Add $x_{n}$ to line $(x_{i},x_{j},x_{k})$; Add $x_{n}$ to $V$
19   end if
20   end for
21   Add line to $\mathcal{L}$
22   Compute mean of line points and add to $\mathcal{M}$
23  end if
24  end for
25  end if
26end for
27for pairs of midpoints $(m_{p},m_{q})\in\mathcal{M}$ do
28  Compute intersection point $x^{*}$ of lines
29    corresponding to $m_{p}$ and $m_{q}$
30  Add $x^{*}$ to $R^{1}$ and record support line
31    indices $\{p,q\}$ in $S(x^{*})$
32end for
33Filter points in $R^{1}$ with $\lvert S(x^{*})\rvert\geq 3$
34return $R^{1}$
Complexity. The worst-case time complexity is very similar to DistinSMOTE, i.e., $\mathcal{O}(n^{2}d+n(kr)^{2})$. While there are two additional factors, namely, $\mathcal{O}(nkr)$ for checking neighbors after identifying a collinear triplet and $\mathcal{O}(n^{2})$ for finding line intersections, these are dominated by existing terms and can be ignored. The practical complexity is much lower, and the attack runs in at most three minutes on all main datasets.
Accuracy Analysis. Next, we analyze ReconSMOTE; the following theorems give theoretical guarantees for its reconstruction rule, achieving perfect precision and a lower bound on expected recall.
Theorem 2 (ReconSMOTE perfect precision).
Under Assumptions 13, the records reconstructed by ReconSMOTE are guaranteed to be real, i.e., the attack achieves perfect precision (with probability 1).
Proof.
By SMOTE construction, each synthetic point lies on a segment strictly between two real endpoints. Accordingly, every detected line in $X^{1}_{syn}$ (obtained from collinear sets identified by the local search in lines 8-17 of Algorithm 3) corresponds uniquely to a pair of real records. The local search correctly groups all synthetic points from that pair into a single collinear set without introducing spurious collinearities.
Under the global non-collinearity assumption, the only point lying on three or more such lines is their shared real endpoint. ReconSMOTE exploits this by intersecting detected lines (lines 23-25) and retaining only points supported by at least three distinct lines (line 27). Since synthetic points lie on exactly one SMOTE line, whereas real endpoints lie on three or more, any retained intersection must be a true real point. Hence, every reconstructed record is real, so precision is $1$ (with probability 1). ∎
For clarity and tractability, the next theorem uses a simplified formulation with directed edges and a Poisson approximation, giving a conservative bound (the approximate bound) that ignores overlapping neighbor relations in the SMOTE graph.
Theorem 3 (ReconSMOTE expected recall (approximate)).
Let $\lambda=\frac{n_{0}-n_{1}}{n_{1}k}$. Under Assumptions 13 and using Poisson approximation (treating the number of synthetic points per segment as Poisson), the expected recall of ReconSMOTE satisfies:
$\mathbb{E}[\mathrm{Recall}]\geq\max\!\left\{0,\ \frac{k\!\left(1-e^{-\lambda}\!\left(1+\lambda+\tfrac{\lambda^{2}}{2}\right)\right)-2}{k-2}\right\}.$
Proof.
At each generation step, SMOTE first selects a minority record $x_{i}$ uniformly from the $n_{1}$ available, and then one of its $k$ nearest neighbors $x_{j}$ uniformly. Thus, each of the $n_{1}k$ possible minority-neighbor (directed) segments is chosen with probability $\frac{1}{n_{1}k}$ at every step.
Let $C_{ij}$ denote the number of synthetic points generated on segment $(x_{i},x_{j})$. Since each of the $n_{0}{-}n_{1}$ synthetic points is assigned independently to a segment with probability $\frac{1}{n_{1}k}$, the vector of all $C_{ij}$ follows a multinomial distribution with $\sum_{ij}C_{ij}=n_{0}{-}n_{1}$, and each $C_{ij}$ is marginally distributed as $\mathrm{Binom}(n_{0}{-}n_{1},\frac{1}{n_{1}k})$. For analytic tractability, we approximate this by $\mathrm{Poisson}(\lambda)$ with mean $\lambda=\frac{n_{0}-n_{1}}{n_{1}k}$. This is standard when $n_{0}{-}n_{1}$ is large and $\frac{1}{n_{1}k}$ is small, which holds in practice.
A segment is reconstructed if $C_{ij}\geq 3$. The probability of this is $p_{\mathrm{edge}}=\Pr\{\mathrm{Poisson}(\lambda)\geq 3\}=1-e^{-\lambda}\!\left(1+\lambda+\tfrac{\lambda^{2}}{2}\right)$. Now consider a fixed record $x_{i}$, and let $S_{i}$ denote the number of reconstructed segments incident to it. Therefore, $\mathbb{E}[S_{i}]=k\,p_{\mathrm{edge}}$. Moreover, by Assumption 3, once $S_{i}\geq 3$, the point $x_{i}$ is uniquely identifiable, since three non-collinear reconstructed segments suffice to triangulate its location. To lower bound $\Pr\{S_{i}\geq 3\}$, observe that $\mathbb{E}[S_{i}]=\mathbb{E}[S_{i}\mathbb{I}\{S_{i}\leq 2\}]+\mathbb{E}[S_{i}\mathbb{I}\{S_{i}\geq 3\}]\leq 2\,\Pr\{S_{i}\leq 2\}+k\,\Pr\{S_{i}\geq 3\}$ (since $S_{i}\leq k$). As $\Pr\{S_{i}\leq 2\}=1-\Pr\{S_{i}\geq 3\}$, this yields $\mathbb{E}[S_{i}]\leq 2+(k-2)\Pr\{S_{i}\geq 3\}$, hence $\Pr\{S_{i}\geq 3\}\ \geq\ \frac{\mathbb{E}[S_{i}]-2}{k-2}=\frac{k\,p_{\mathrm{edge}}-2}{k-2}:=A_{id}$.
This is the probability that $x_{i}$ is identifiable. Since recall is the fraction of minority records that are identifiable, its expectation equals the average of these probabilities over all $n_{1}$ records. Because each $x_{i}$ is treated symmetrically in SMOTE and we look at directed segments, the average equals the bound derived above. Hence we obtain the stated lower bound on the expected recall. ∎
Remarks. By rearranging Equation 3, we get $1-\mathbb{E}[\mathrm{Recall}]\;\leq\;\frac{k}{k-2}\,e^{-\lambda}(1+\lambda+\tfrac{\lambda^{2}}{2})$, which in turn means $\mathbb{E}[\mathrm{Recall}]\to 1$ as $\lambda\to\infty$, with a convergence rate exponential in $\lambda(=\frac{n_{0}-n_{1}}{n_{1}k}=\frac{r-1}{k})$.
In Appendix A, we provide a more detailed analysis and derive a tighter bound without the simplifications of Theorem 3; we call it the exact bound. Finally, in Appendix B, we visualize the differences between the bounds under various conditions.
5 Experimental Evaluation
We now evaluate the effectiveness of our novel attacks, along with additional methods geared to assess privacy leakage in SMOTE, both as a data augmentation and synthetic data generation technique. Specifically, we consider: 1) current practices such as naive distinguish (via a classifier) and privacy metrics (i.e., DCR from synthetic to real records [65]), 2) state-of-the-art Membership Inference Attacks (MIAs) [56, 6], which to the best of our knowledge have not yet been applied against SMOTE, and 3) the DistinSMOTE and ReconSMOTE attacks. Overall results are summarized in Table 1.
[…]
ecoli imU 8.6 336 7
abalone 7 9.7 4,177 10
car_eval_34 vgood 12 1,728 21
solar_flare_m0 M-0 19 1,389 32
car_eval_4 vgood 26 1,728 21
yeast_me2 ME2 28 1,484 8
mammography minority 42 11,183 6
abalone_19 19 130 4,177 10
Figure 2: Main datasets overview, where $r$ denotes the imbalance ratio ($n_{0}/n_{1}$), $n$ the number of records, and $d$ the number of features.
Datasets. We conduct our main experiments on eight standard imbalanced datasets, each with a binary classification task, obtained from the imblearn library [40] (originally from the UCI ML Repository) and used in prior work [13, 55] These datasets vary significantly in size (336 to 11,183 records), dimensionality (6 to 32 features), imbalance ratios (8.6 to 130), and prediction task (target), as shown in Table 2.
Implementations. We use the standard imblearn [40] implementation of SMOTE and sklearn [51] for classifiers. Both DistinSMOTE and ReconSMOTE are highly efficient, running in under three minutes on any dataset from Table 2 on an Apple M4 MacBook with 24GB RAM. The naive methods are similarly fast, while the MIAs take up to 30 minutes per dataset. We will release the source code for our attacks along with the final version of the paper.
5.1 Augmented Data
We compare the three approaches on augmented data, with results for all datasets shown in Table 2.
Naive Distinguish is a popular but arguably misguided approach for telling apart real and synthetic records by training a classifier [58, 19, 53, 11]. Half of the real and half of the synthetic data are used to train a Random Forest classifier, with testing performed on the remaining data. For each dataset, we run 5 independent SMOTE generations and train 5 classifiers per run, reporting averaged results. The method severely underestimates privacy risk (see the two leftmost columns in Table 2; precision and recall $\approx 0$) as it is capable of capturing only distributional differences, not record-level leakage.
MIA. Next, we evaluate MIAs [56, 6] using the repeated classification game from Section 2. For a given target record, we train 200 classifiers (a multi-layer perceptron with two hidden layers) on augmented datasets generated via SMOTE: half of the training datasets include the target record, and half exclude it. We then use the classifiers’ predictions on the target to simulate an adversary’s confidence in distinguishing membership, and calculate AUC. Following prior work [62, 27], we train target-specific attacks in a leave-one-out setting, which provides a more accurate estimate of privacy leakage. This procedure is repeated for 100 randomly selected targets (or all minority records), and we report the average. Overall, this requires training roughly 20k SMOTE models and classifiers per dataset.
Looking at Table 2 (fourth column), the average AUC is 0.68, with half of the datasets exceeding 0.7, which indicates substantial privacy leakage. The lowest scores appear in datasets with the smallest imbalance (ecoli and abalone), where the proportion of synthetic data is relatively low. Mammography also shows a low score, likely because its large number of records reduces the influence of any single individual. These results are therefore not entirely surprising.
We also conduct an additional MIA experiment, training classifiers solely on the real data, to test the intuition that SMOTE enhances the sensitivity of minority records in the augmented data, as they directly contribute to generating synthetic samples. As expected, targets become more vulnerable when augmentation is applied – average AUC increases by 17% (comparing the third and fourth columns in Table 2). Larger imbalance further amplifies this effect. While similar intuitions have been noted previously [55], they were not supported by empirical evidence.
[…]
$D_{aug}$ $D_{real}$ $D_{aug}$ $D_{aug}$
(Precision) (Recall) (AUC) (AUC) (Precision) (Recall)
ecoli 8.6 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.50 $\pm$ 0.04 0.50 $\pm$ 0.05 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
abalone 9.7 0.03 $\pm$ 0.03 0.00 $\pm$ 0.01 0.57 $\pm$ 0.03 0.58 $\pm$ 0.04 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
car_eval_34 12 0.01 $\pm$ 0.02 0.01 $\pm$ 0.01 0.60 $\pm$ 0.03 0.73 $\pm$ 0.08 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
solar_flare_m0 19 0.01 $\pm$ 0.03 0.00 $\pm$ 0.01 0.79 $\pm$ 0.03 0.97 $\pm$ 0.03 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
car_eval_4 26 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.59 $\pm$ 0.03 0.75 $\pm$ 0.10 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
yeast_me2 28 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.51 $\pm$ 0.04 0.57 $\pm$ 0.09 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
mammography 42 0.01 $\pm$ 0.02 0.00 $\pm$ 0.00 0.54 $\pm$ 0.03 0.56 $\pm$ 0.04 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00
abalone_19 130 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.58 $\pm$ 0.05 0.80 $\pm$ 0.12 0.99 $\pm$ 0.02 1.00 $\pm$ 0.00
average 0.01 $\pm$ 0.01 0.00 $\pm$ 0.00 0.58 $\pm$ 0.03 0.68 $\pm$ 0.07 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
Table 2: Privacy attacks vs. augmented data.
DistinSMOTE. Finally, we run DistinSMOTE on 25 SMOTE generations and report average precision/recall (two rightmost columns in Table 2). As expected from our analysis, we achieve perfect results across all datasets and imbalance levels. This shows that merely knowing SMOTE was used for augmentation is enough for an adversary to perfectly identify real records with minimal effort.
5.2 Synthetic Data
Next, we evaluate all attacks on synthetic data; see Table 3.
[…]
(Accuracy) (AUC) (Precision) (Recall)
ecoli 8.6 0.19 $\pm$ 0.15 0.93 $\pm$ 0.05 1.00 $\pm$ 0.00 0.43 $\pm$ 0.02
abalone 9.7 0.21 $\pm$ 0.17 0.65 $\pm$ 0.07 1.00 $\pm$ 0.00 0.62 $\pm$ 0.01
car_eval_34 12 0.00 $\pm$ 0.00 0.97 $\pm$ 0.01 1.00 $\pm$ 0.00 0.83 $\pm$ 0.03
solar_flare_m0 19 0.03 $\pm$ 0.06 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.95 $\pm$ 0.02
car_eval_4 26 0.01 $\pm$ 0.04 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
yeast_me2 28 0.20 $\pm$ 0.12 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
mammography 42 0.25 $\pm$ 0.15 0.91 $\pm$ 0.04 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
abalone_19 130 0.37 $\pm$ 0.14 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
average 0.16 $\pm$ 0.10 0.93 $\pm$ 0.02 1.00 $\pm$ 0.00 0.85 $\pm$ 0.01
Table 3: Privacy attacks vs. synthetic data.
Naive Metrics. A widely used approach for evaluating privacy in synthetic data is the Distance to Closest Record (DCR) [65], which measures the average distance between synthetic and real records. DCR has been commonly applied to SMOTE and modern diffusion models [37, 63, 50, 45], but its interpretation is limited – an average distance alone provides little insight into privacy risks. To address this, we use a linkability attack [25], which builds on DCR and reports the accuracy with which an adversary could link two partial feature sets of a real record using synthetic data. For each dataset, we train 5 SMOTE models and evaluate linkability 5 times with varying feature subsets.
The results are unstable (see the leftmost column in Table 3): scores differ from zero only for low-dimensional settings ($d\leq 10$), while higher-dimensional datasets yield large variances that render DCR unreliable. This is expected, as DCR treats all features equally and is known to be an inadequate privacy measure [4, 22, 61].
MIA. We evaluate MIAs on synthetic data using the repeated classification game (similar to Section 5.1). We rely on the GroundHog attack [59], one of the most popular MIAs for synthetic tabular data. GroundHog extracts statistical features from generated datasets – such as column-wise minimum, mean, median, maximum, and pairwise correlations – and uses them to train a meta-classifier, which is then applied to unlabeled real and synthetic feature sets. To generate training features, we train 400 SMOTE models for in/out training features and another 200 SMOTE models for in/out testing features. Repeating this for 100 targets yields about 60k models per dataset.
Figure 3: ReconSMOTE recall and lower bounds (per dataset).
Figure 4: ReconSMOTE performance w/ varying $r$ (all datasets).
As shown in Table 3 (second column), this results in substantial privacy leakage: AUC exceeds 0.9 in all but one dataset. The exception is abalone, which has the second-lowest imbalance ratio and the second-largest number of records, potentially leading to lower sensitivity. When imbalance increases (abalone_19), the MIA AUC rises to 1. Overall, these results demonstrate that SMOTE-generated data is highly susceptible to MIAs, even beyond trivial cases where data domain characteristics mainly drive leakage [23, 24].
ReconSMOTE. Next, we apply ReconSMOTE on 25 SMOTE generations per dataset, reporting average precision/recall (last two columns in Table 3). The attack achieves perfect precision on all datasets, which, as motivated in Section 4, is the most critical metric for reconstruction. Recall is also very high (see Figure 3), with an average of 0.85. It increases quickly with class imbalance, reaching 1 when $r\geq 20$. The recall values (per dataset) are in line with the expected approximate/exact bounds predicted by Theorem 3 and 4.
To further validate the expected bounds at finer granularity, we vary the imbalance ratio $\{5,10,20,25,50,75,100\}$ across all datasets and plot the average performance in Figure 4. As expected, recall increases exponentially with $r$ (for fixed $k$), reaching 1 around imbalance 20. Overall, these findings highlight the risks of relying on SMOTE for synthetic data generation: in realistic settings, an adversary can reconstruct all real records with perfect confidence.
5.3 Take-Aways
We show that MIAs achieve high AUC across numerous targets vs. SMOTE: 0.68 against classifiers trained on augmented data and 0.93 against synthetic data. Moreover, the sensitivity of minority records increases by an average of 17% when classifiers are trained on augmented rather than original training data. Finally, our attacks, DistinSMOTE and ReconSMOTE, are able to i) distinguish real minority records from synthetic ones in augmented data, and ii) reconstruct real minority records from synthetic data with minimal assumptions and near-perfect accuracy.
6 DistinSMOTE and ReconSMOTE with Relaxed Assumptions
In this section, we test the robustness of our attacks, DistinSMOTE and ReconSMOTE, on augmented/synthetic data while relaxing our assumptions one at a time, e.g., using high-dimensional data, mixed-type data, and $k=2$; results are shown in Table 45, and 6 in Appendix C. Additionally, in Appendix C, we evaluate our attacks on perturbed SMOTE datasets (i.e., linear interpolation with added random noise) and provide a heuristic for running the attacks without knowledge of $k$ and $r$.
High-Dimensional Data. First, we test how the attacks scale to two high-dimensional datasets – up to 96,690 records and 50 features (see Table 4 and Appendix C for results and details about the higgs and miniboone datasets, both with $r=25$). Both DistinSMOTE and ReconSMOTE achieve perfect precision and recall and, as analyzed in Section 4, scale well with $n$ and $d$, taking no more than 8 minutes per dataset, which is highly practical.
Mixed-Type Data. Next, we show that our attacks also apply to mixed-type data, relaxing Assumption 1; see Table 5 and Appendix C for details on the cardio and churn datasets, each containing an equal mix of numerical/categorical features, and with $r=25$. To generate data, we use SMOTE-NC [7] – introduced in the original SMOTE paper and designed for mixed data – via the standard imblearn implementation [40]. For the attacks, we simply ignore categorical features and operate on the continuous ones. As before, we obtain perfect precision and very high recall on both datasets. This approach can also be applied to one-hot encoded data.
SMOTE with $\mathbf{k=2.}$ Finally, we evaluate the attacks on the eight main datasets using SMOTE with $k=2$, relaxing Assumption 3 (see Table 6 in Appendix C). The performance of DistinSMOTE remains unaffected, achieving perfect precision and recall, as expected. As for ReconSMOTE, average recall drops to 0.52 (a 39% decrease compared to SMOTE with $k=5$ in Table 3) because each real record participates in fewer lines, making it harder to reach the support threshold; only records that serve as neighbors to other records beyond their two closest neighbors can be successfully reconstructed. Precision remains perfect as all reconstructed points are still accurate. Nevertheless, reconstructing half the real minority records with perfect confidence is a serious privacy breach.
7 Conclusion
Our work highlights the fundamental privacy limitations of SMOTE [7], one of the most widely adopted techniques for improved learning on imbalanced data. The effectiveness of our novel, near assumption-free attacks (DistinSMOTE and ReconSMOTE), demonstrates that real minority records – precisely the ones SMOTE aims to better represent – are exposed to significant, previously underestimated privacy risk. Importantly, this also shows that using SMOTE as a baseline with DCR to evaluate privacy is unreliable and can provide a false sense of security. Nonetheless, SMOTE remains an effective and easy-to-use technique in non-privacy-sensitive applications where utility is the primary concern. We are confident our findings will be valuable to researchers and practitioners deploying solutions that process or release sensitive data, motivate them to avoid SMOTE as a privacy benchmark, and encourage them to adopt more robust privacy-preserving techniques.
Limitations and Future Work. Our attacks currently operate on continuous data and are primarily tested on the original SMOTE implementation. While certain numerical instabilities/edge cases are theoretically possible (e.g., a synthetic point appearing collinear with two unrelated real points), their probability is effectively zero in high-dimensional datasets with high numerical precision, and we did not observe any such case in our experiments. Additionally, our findings generalize to many SMOTE variants – such as BorderlineSMOTE [28], ADASYN [30], SVMSMOTE [47], and cluster/hybrid-based methods [16] – as they all rely on line-segment interpolation to generate synthetic samples. In contrast, our attacks are unlikely to be successful against variants like G-SMOTE [15] and GI-SMOTE [10], which generate synthetic points within regions rather than strictly along lines. Nevertheless, these variants are not inherently privacy-preserving and are still likely to remain vulnerable to MIAs. Extending our attacks and developing robust defenses for such methods is a promising direction for future work.
Privacy-preserving variants of SMOTE have also been proposed under the framework of Differential Privacy [17, 18], including DP-SMOTE [41], which adds noise when estimating point distributions/nearest neighbors, and SMOTE-DP [66], which combines SMOTE with a DP generative model. However, SMOTE-DP largely ignores SMOTE’s increased sensitivity of minority records [39, 41, 55], a gap we confirm empirically (see Section 5.1). As none of these approaches provides open-source implementations, we leave evaluating their effectiveness to future work.
Ethics Statement. Our work does not involve attacking live systems or private datasets. Our goal is to demonstrate the importance of emphasizing privacy considerations and relying on established notions of privacy when processing sensitive, imbalanced data in critical domains. To make our work reproducible, we release the source code for our attacks and experiments: https://github.com/sascommunities/smote-mirrors.
Appendix A Tighter Lower Bound on ReconSMOTE Recall
In this section, we derive a tighter lower bound on the recall of ReconSMOTE by relaxing two of the assumptions in Section 4.3, namely the Poisson approximation and the one-directional counting of $C_{ij}$. Specifically, we use the exact Binomial distributions and count $C_{ij}$ in both directions to obtain more accurate values.
Recall that at each step of the SMOTE algorithm, we choose an $x_{i}\in X^{1}_{real}$ uniformly at random and then independently select one of its $k$ nearest neighbors uniformly at random. To capture this structure, we represent the minority data by a KNN graph $G=(X^{1}_{\text{real}},E)$, where edges $E$ represent the neighboring relations among the minority samples. As before, we use $N(x_{i})$ to denote the $k$ nearest points to $x_{i}$ from the minority set. For each $x_{i}\in X^{1}_{real}$, we add an edge $E_{i\to j}$ whenever $x_{j}\in N(x_{i})$. Each synthetic data generated by SMOTE is associated with exactly one edge. Let $\alpha$ denote the probability that a nearest-neighbor relation is mutual; i.e., the probability that if $x_{j}\in N(x_{i})$ then also $x_{i}\in N(x_{j})$. In this case, a synthetic point lies on $E_{i\to j}$ if it was generated along $E_{i\to j}$ or along $E_{j\to i}$. If $\alpha=1$, then all nearest-neighbor edges are mutual, and $\alpha=0$ corresponds to a completely one-sided nearest-neighbor graph, for which we usually refer to those edges as exclusive.
Recall that an edge is reconstructed if at least three synthetic records lie on its segment, and a real $x_{i}$ is identifiable if there exist three reconstructed edges incident to $x_{i}$. We denote the number of synthetic data points generated between $x_{i}$ and $x_{j}$ by $C_{ij}$. For generating each synthetic point, SMOTE performs these selections independently, assigning every new sample to exactly one of the $n_{1}k$ directed edges with equal probability $1/(n_{1}k)$. This independence is not an additional assumption – it follows directly from the random sampling mechanism of SMOTE. Consequently, the vector of all $C_{ij}$ follows a multinomial distribution with $\sum_{ij}C_{ij}=n_{0}{-}n_{1}$, and each component $C_{ij}$ is marginally $\mathrm{Binom}(n_{0}{-}n_{1},\frac{1}{n_{1}k})$.
Because some directed edges in the SMOTE KNN graph represent mutual neighbor relationships, certain edges overlap. We address this by distinguishing between one-way and mutual edges. Let $B_{ij}:=\mathbb{I}\{x_{i}\in N(x_{j})\;\mathrm{and}\;x_{j}\in N(x_{i})\}$ (mutuality indicator) with $\Pr\{B_{ij}=1)=\alpha$. Then
$C_{ij}\,\big|\,B_{ij}=0\sim\mathrm{Binom}\!\Big(n_{0}{-}n_{1},\tfrac{1}{n_{1}k}\Big),\qquad C_{ij}\,\big|\,B_{ij}=1\sim\mathrm{Binom}\!\Big(n_{0}{-}n_{1},\tfrac{2}{n_{1}k}\Big).$
Consider the following assumption regarding the structure of neighboring relations around each real minority record.
Assumption 4 (Local non-degeneracy).
For any $x_{i}\in X^{1}_{\text{real}}$, all edges $\{E_{i\to j}:x_{j}\in N(x_{i})\}$ have pairwise distinct directions.
In the analysis of this section, we use Assumption 4 in place of Assumption 2 (global non-collinearity), as it is a weaker, localized condition sufficient for establishing the lower bound on reconstruction recall. In particular, if Assumption 2 holds, then Assumption 4 automatically follows. Under Assumption 4, the intersection of any three reconstructed edges incident to $x_{i}$ uniquely identifies $x_{i}$.
Lemma 1 (Reconstructed edge probability).
For any edge of $G$,
$\Pr\{C_{ij}\geq 3\}{=}(1-\alpha)\Pr\{\mathrm{Binom}(n_{0}{-}n_{1},\tfrac{1}{n_{1}k})\geq 3\}+\alpha\Pr\{\mathrm{Binom}(n_{0}{-}n_{1},\tfrac{2}{n_{1}k})\geq 3\}{=:}p_{\mathrm{edge}}(\alpha).$
Proof.
Condition on $B_{ij}$ and compute the average. If $B_{ij}=0$, then only one direction contributes to the count; if $B_{ij}=1$, both directions do. ∎
Lemma 2 (Lower-bound on per-node identifiability).
Fix $x_{i}\in X^{1}_{\text{real}}$ and its $k$ outgoing directed edges $\{E_{i\to j}:x_{j}\in N(x_{i})\}$. Then, we have
$\Pr\{x_{i}\ \mathrm{identifiable}\}\;\geq\;\max\!\left\{\,0,\ \frac{k\,p_{\mathrm{edge}}(\alpha)-2}{k-2}\right\}=:L_{id},$ (2)
and this lower bound is tight.
Proof.
Declare an edge reconstructed if $C_{ij}\geq 3$ and set $E_{i\to j}^{\mathrm{rec}}:=\mathbb{I}\{C_{ij}\geq 3\}$. Let $S_{i}=\sum_{j=1}^{k}E_{i\to j}^{\mathrm{rec}}$. From Lemma 1, each edge has marginal $\Pr\{E_{i\to j}^{\mathrm{rec}}=1\}=p_{\mathrm{edge}}(\alpha)$, so $\mathbb{E}[S_{i}]=k\,p_{\mathrm{edge}}(\alpha)$. Then, we have
$\mathbb{E}[S_{i}]$ $=\mathbb{E}\!\big[S_{i}\,\mathbb{I}\{S_{i}\leq 2\}\big]+\mathbb{E}\!\big[S_{i}\,\mathbb{I}\{S_{i}\geq 3\}\big]$
$\leq\mathbb{E}\!\big[2\,\mathbb{I}\{S_{i}\leq 2\}\big]+\mathbb{E}\!\big[k\,\mathbb{I}\{S_{i}\geq 3\}\big]\qquad\text{(since $S_{i}\leq k$ a.s.)}$
$=2\,\Pr\{S_{i}\leq 2\}+k\,\Pr\{S_{i}\geq 3\}.$
Since $\Pr\{S_{i}\leq 2\}=1-\Pr\{S_{i}\geq 3\}$, we obtain
$\mathbb{E}[S_{i}]\ \leq\ 2+(k-2)\,\Pr\{S_{i}\geq 3\},$
hence
$\Pr\{S_{i}\geq 3\}\ \geq\ \frac{\mathbb{E}[S_{i}]-2}{k-2}=\frac{k\,p_{\mathrm{edge}}(\alpha)-2}{k-2}.$
Truncating at 0 accommodates the trivial case $k\,p_{\mathrm{edge}}(\alpha)\leq 2$.
Moreover, the bound cannot be improved using only the edge-wise success probabilities. Consider constructing $(E_{i\to j}^{\mathrm{rec}})_{j=1}^{k}$ so that $S_{i}=\sum_{j=1}^{k}E_{i\to j}^{\mathrm{rec}}$ takes values only in $\{2,k\}$. Choose the mixture weights so that $\mathbb{E}[S_{i}]=k\,p_{\mathrm{edge}}(\alpha)$. In this case, the inequality holds with equality, meaning that the lower bound is tight. ∎
Theorem 4 (ReconSMOTE expected recall (exact)).
Under Assumptions 13, and 4, we have
$\mathbb{E}[\mathrm{Recall}]\coloneqq\mathbb{E}[\frac{1}{n_{1}}\#\{x_{i}\in X^{1}_{\text{real}}:\ x_{i}\ \mathrm{identifiable}\}]\;\geq\;L_{id},$ (3)
where $L_{id}$ is defined in Equation 2.
Proof.
By Lemma 2, we have $\Pr\{x_{i}\ \mathrm{identifiable}\}\geq L_{id}$ for every $i$, so
$\mathbb{E}[\mathrm{Recall}]=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\Pr\{x_{i}\ \mathrm{identifiable}\}\ \geq\ L_{id}.$
Appendix B Lower Bounds of ReconSMOTE Recall Visualizations
To complement the theoretical results in Section 4.3 (approximate bound, $A_{id}$) and Appendix A (exact bound, $L_{id}$), we visualize the bounds under different conditions.
We start by showing the ratio between the exact bound and the approximate bound as a function of the imbalance ratio $r$ in Figure 5. When $\alpha=0$ (Figure 5(a)), the approximate bound closely matches the exact one for all $k$, especially for larger imbalance ratio. In contrast, even a small amount of mutuality ($\alpha=0.1$; Figure 5(b)) introduces a noticeable deviation of the lower bound.
(a) $\alpha=0$
(b) $\alpha=0.1$
Figure 5: Exact/approximate bound ratios for two levels of mutuality.
(a) Lower bound $L_{id}$ vs. $r$
(b) Heatmap of $L_{id}$ over $(r,k)$
Figure 6: Comparison of two visualizations of the exact bound.
Figure 7: Exact bound as a function of $\alpha$ for different imbalance ratios $r$.
Next, in Figure 6, we focus on the exact bound $L_{id}$ under varying imbalance ratios $r$ and neighborhood sizes $k$ (with $n_{1}=100$ and $\alpha=0.5$ fixed). Figure 6(a) shows that $L_{id}$ steadily increases as the oversampling ratio $r$ rises. For small $k$, even a moderate oversampling ratio results in significant identifiability. However, for larger $k$, a higher oversampling ratio is required. The heatmap in Figure 6(b) clearly illustrates this interaction. In the upper-left area, where $k$ is small and $r$ is large, $L_{id}$ quickly approaches 1. This indicates almost perfect identifiability. In contrast, in the lower-right area, where $k$ is large and $r$ is small, $L_{id}$ is close to zero. This suggests that the reconstructed edges are not dense enough to reach high identifiability. Overall, these plots confirm the trade-off: identifiability improves with oversampling, but its efficiency depends strongly on the neighborhood parameter $k$.
Finally, Figure 7 presents the exact bound $L_{id}$ as a function of $\alpha$ for several imbalance ratios $r$. The curves illustrate the sensitivity of $L_{id}$ to the graph structure. For example, when $r=10$, small increases in $\alpha$ would lead to substantial changes in $L_{id}$, highlighting how mutuality in the KNN graph strongly influences privacy leakage.
Appendix C DistinSMOTE and ReconSMOTE with Relaxed Assumptions
In this section, we present results for our attacks under relaxed assumptions – on high-dimensional data (Table 4), mixed-type data (Table 5), SMOTE with $k=2$ (Table 6), and perturbed data (Table 7 and 8). For the first two experiments, we use datasets different from the eight main datasets in Table 2. Namely, we use higgs and miniboone from OpenML [60] as high-dimensional data, and cardio and churn from Kaggle as mixed-type data. These datasets are used in relevant prior work [37]. For all datasets, the minority class is undersampled so the imbalance is 25. The results of the first three experiments are discussed in Section 6.
Next, we discuss the results of the forth experiment – running our attacks on perturbed SMOTE data.
[…]
(Precision) (Recall) (Precision) (Recall)
higgs 25 47,976 28 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
miniboone 25 96,690 50 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
Table 4: Privacy attacks vs. augmented/synthetic data with high-dimensional data.
[…]
(num, cat) (Precision) (Recall) (Precision) (Recall)
cardio 25 7,256 (5, 6) 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00
churn 25 8,269 (5, 5) 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.01
Table 5: Privacy attacks vs. augmented/synthetic data with mixed-type data.
[…]
(Precision) (Recall) (Precision) (Recall)
ecoli 8.6 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.43 $\pm$ 0.06
abalone 9.7 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.42 $\pm$ 0.02
car_eval_34 12 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.61 $\pm$ 0.02
solar_flare_m0 19 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.40 $\pm$ 0.02
car_eval_4 26 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.60 $\pm$ 0.00
yeast_me2 28 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.58 $\pm$ 0.02
mammography 42 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.61 $\pm$ 0.01
abalone_19 130 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.49 $\pm$ 0.01
average 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.52 $\pm$ 0.02
Table 6: Privacy attacks vs. augmented/synthetic data by SMOTE with $k=2$.
SMOTE with Perturbed Linear Interpolation. We test the robustness of our attacks on perturbed data by adding column-wise noise in the range $10^{-10},10^{-7},10^{-5},10^{-3}$}, ensuring that no synthetic record lies exactly on the line between its generating real records (see Table 7 and 8). Similarly, we use tolerance levels in the same range for detecting lines/intersections within our attacks. We use the yeast_me2 dataset from our main experiments (Table 2).
Perhaps surprisingly, both DistinSMOTE and ReconSMOTE remain highly effective when the added noise is small ($\leq 10^{-7}$), achieving near-perfect performance. This shows that our attacks can generalize to SMOTE variants that use non-strictly linear interpolation. However, the performance of both attacks drops sharply when larger noise is injected – though such noise levels would likely degrade downstream utility as well. Across all noise settings, we observe a consistent trend: for each noise level, both attacks achieve their best precision (the more important metric, as already discussed) when the tolerance parameter matches the injected noise level.
DistinSMOTE SMOTE augmented data w/ noise per column (Prec./Rec.)
noise $10^{-10}$ noise $10^{-7}$ noise $10^{-5}$ noise $10^{-3}$
tol. $10^{-10}$ 0.96 1.00 0.88 1.00 0.04 1.00 0.04 1.00
tol. $10^{-7}$ 0.91 0.99 0.91 0.99 0.07 1.00 0.04 1.00
tol. $10^{-5}$ 0.41 0.80 0.37 0.78 0.07 0.93 0.04 0.92
tol. $10^{-3}$ 0.09 0.31 0.09 0.31 0.06 0.82 0.06 0.80
Table 7: DistinSMOTE vs. perturbed augmented data, on yeast_me2.
ReconSMOTE SMOTE synthetic data w/ noise per column (Prec./Rec.)
noise $10^{-10}$ noise $10^{-7}$ noise $10^{-5}$ noise $10^{-3}$
tol. $^{-10} 1.00 1$.00 0.86 0.96 0.00 0.00 0.00 0.00
tol. $^{-7} 1.00 1$.00 0.96 1.00 0.00 0.00 0.00 0.00
tol. $10^{-5}$ 0.99 0.97 0.85 0.84 0.71 0.10 0.00 0.00
tol. $10^{-3}$ 0.63 0.61 0.55 0.53 0.10 0.09 0.00 0.00
Table 8: ReconSMOTE vs. perturbed synthetic data, on yeast_me2.
Finally, we relax another assumption – running our attacks without prior knowledge of $k$ and $r$.
SMOTE with Unknown $k$ and $r$. We describe heuristics for running DistinSMOTE and ReconSMOTE without knowing SMOTE’s parameters $k$ and $r$. In practice, precise estimates are unnecessary: the neighborhood search only needs to be wide enough to include the real records that generated a given (synthetic) record. Overestimating $k$ or $r$ does not reduce precision/recall, it only increases runtime. Thus, a simple strategy would be to skip parameter estimation altogether and simply use a large neighborhood search (e.g., 10-25% of the dataset).
To estimate $k$ or $r$ more accurately, the adversary can reuse the same sub-procedures employed in the attacks. For a given record, a large neighborhood search (e.g., 10-25% of the dataset) identifies all neighbors on the same line. For augmented data, the adversary can locate an endpoint (a real record), run a second search around it, and infer: i) the number of lines pointing to this record (an overestimate of $k$), and ii) the number of records per line (a rough approximation of $r/k$). For synthetic data, the adversary can instead detect three intersecting lines around the record, identify their intersection (a real record), and proceed as above. Repeating this procedure and averaging yields stable estimates.